# Quasiparticle states of hexagonal BN: A van der Waals density functional study

### v1

@@ II.2 CX in Koopmans DFT
From a theoretical perspective, the eigenstates of KS-DFT, obtained using a regular XC functional, provide a reasonable first approximation for studying charge excitations in [hohenberg1964, kohn1965]. Still, while KS-DFT serves us to accurately predict GS properties, its eigenvalues $\epsilon_{i}^{\text{KS}}$ and orbitals $\phi_{i}^{\text{KS}}(\mathbf{r})$ do not directly describe charged excitations such as the set of vertical ionization energies, i.e., the QPs. This is true even though that all the necessary information to describe a frequency moment of the fully interacting spectral function is, in fact, available by simply combing terms of the KS-DFT energy descriptions with also an evaluation of the kinetic-correlation energy [Burke97, signatures, ChiDFT23]. Even a PBE-based evaluation of those DFT terms yields a value for the moment of the interacting spectral function that closely reflects a full multi-reference configuration integral studies for small molecules [ChiDFT23].
The limitation of using KS-DFT orbitals as a predictor for QPs arises because KS-DFT relies on using (regular, nonhybrid) XC functionals to define an effective local potential contribution [helujpc1971, kronik2012] to set the orbital structure. Plotting the total energy as a function of the number of electrons then yields a convex curvature [kronik2012]. Use of regular XC functionals to thus directly set KS-DFT orbitals implies a lack of the PWL that is, for example, needed for a DFT-type framework to also satisfy a thermodynamics constraints on charge transfers [perdew1982, kronik2012].
Additional information, concealed as MBPT input to the XC functional designs, is needed to also recover accurate results for QP levels [davo2010, kronik2012, WiOhHa21]. Fortunately the information is still formally retained in the MBPT input and physics logic of good XC designs [ChiDFT23]. By switching to Koopmans-compliant DFT and using the KI-DFT framework with a linear-response extensions of an underlying XC functional (like PBE or the CX version of the vdW-DF method), we can leverage more of the underlying MBPT contents [helujpc1971, hyldgaard2020, ChiDFT23] of such robust XC designs also for accurate QP predictions [colonna2022, schubert2023, OTRSHalga, OTRSHgap, WiOhHa21, JPCM2025].
On a formal level, KI-DFT employs a class of orbital-density-dependent functionals (ODDFs) designed to enforce generalized PWL. Specifically, the ODDFs are crafted to enforce that we have a linear energy change with respect to variations in the electron occupation [perdew1982, kronik2012, davo2010]. The CX-based adjustments for KI-DFT, what can also be called the Koopmans functional, is given:
@@
Here $\{\alpha_{i}\}$ are screening parameters that account for electronic relaxation effects and $\rho_{i}$ are orbital-specific density variations $\rho_{i}(\mathbf{r})=f_{i}|\phi_{i}(\mathbf{r})|^{2}$; with $0<f_{i}<1$ [davo2010]. We are using the KI form [borghi2014, nguyen2018, colonna2022, linscott2023] so that the formal KI-functional specification, $E^{\text{KI-CX}}_{\text{xc}}\left[\rho,\{\rho_{i}\}\right]$ reduces exactly to the underlying CX functional form when there is an integer charge in the unit cell (and hence KI-CX does not change KS-DFT structure predictions obtained by CX). However, for partial orbital occupation, the Koopmans-correction term $\Pi_{i}^{\text{CX}}$ in Eq. (7) restores PWL: It converts the nonlinear dependence of the CX functional on changes in orbital-specific occupations, $\int\rho_{i}(\mathbf{r})d\mathbf{r}$, into a linear form between any two integer occupations. This approach suffices to convert the KS-DFT input (i.e., the choice of the XC) into the KI-DFT tool for making corresponding QP predictions [nguyen2018, gennaro2022, linscott2023]. Further details are provided in Refs. [davo2010, linscott2023].
@@
Although this method rigorously provides both variational orbitals and $\{\alpha_{i}\}$, its application remains challenging due to the requirement of multiple constrained DFT and Koopmans calculations in supercells for periodic systems. To address this limitation, the kcw formulation, discussed in Refs. [baroni2001, colonna2022, gennaro2022], employs a second-order Taylor expansion of $\Pi_{i}[\rho,\rho_{i}]$ and use of DFPT in primitive cells. This approach assumes coincidence between variational orbitals and maximally localized Wannier functions (MLWFs) $\quantity{w_{i\mathbf{k}}}$) [mazari2012], that are computed by use of the same underlying functional. These MLWFs simplify the minimization of $E^{\text{KI-CX}}_{\text{xc}}\left[\rho,\{\rho_{i}\}\right]$.
@@
where $\mathcal{V}^{\text{KI}}_{i\mathbf{0}}$ is the orbital density dependent potential (ODDP) and corresponds to the variation $\delta\sum_{j}\Pi_{j}/\delta\rho_{i}$, and $\hat{h}^{\text{CX}}$ is the effective potential defined by CX functional [burke]. The subscript index $\mathbf{0}$ in $\mathcal{V}^{\mathrm{CX}}_{\mathbf{0}i}$ and $\alpha_{\mathbf{0}i}$ refers to a decomposition into a set of independent problems, where a supercell calculation with $\Gamma$-point sampling corresponds to sampling the folded Brillouin zone of the primitive cell [colonna2022]. Hamiltonian diagonalization in Eq. (8),
@@
yields so-called canonical orbitals that corresponds to the eigenvectors $\tilde{w}_{\mathbf{k}i}$. The eigenvalues of Eq. (9), $\varepsilon_{\mathbf{k}i}$ define the KI-CX (or KI-DFT) predictions for the QP energies. We note that the canonical orbitals have a Block-wavefunction-like character; They are delocalized as they are linear combination of MLWFs $w_{\mathbf{k}i}$. It has been shown that use of an ODDP can provide a QP approximation to the frequency-dependent contraction of the many-body self-energy term [ferretti2014]. The implication is that ODDFs have the flexibility to accurately describe both total energies and QP excitations.
@@
Appendix A provides details on a formal representation of a plane-wave and grid-based determination of Eq. (10) using the Roman-Soler spline formulation [roso09] for (the functional nature [dion2004, thonhauser2007, hybesc14] of) the nonlocal-correlation kernel $\Phi_{c}^{\text{nl}}[\rho](\mathbf{r},\mathbf{r^{\prime}})$ entering Eq. (6). The formal expressions are already coded in the linear-response module (kcw) of QE thanks to a previous study documenting fingerprints of vdW and nonlocal-correlation interactions on phonon predictions [sabatini2016].
@@ III Computational details
We compute the QP states by the Koopmans-DFT code package [koopweb2025, linscott2023], using the KI-DFT versions, and the above-summarized kcw formulation [colonna2022]. All those calculations as well corresponding KS-DFT studies are done using norm-conserving PBE-type (scalar-relativistic) pseudopotentials from the PseudoDojo library [vansetten2018]. That is, we employ dual projectors and generalized norm conservation to accurately reproduce all-electron potential binding and scattering properties. Appendix B presents a discussion of numerical stability for the QP-energy predictions with the wavefunction energy cut off $E_{\rm cut}^{\rm ef}$ (and, for the strictly two-dimensional BN1 system, the height $c$ of the modeling cell). This analysis leads us to use $E_{\rm cut}^{\rm wf}=200$ Ry (400 Ry) for bulk (single-sheet) h-BN, while doubling the grid for the density variations; Exceptions are explicitly stated.
@@
In the initial step, Koopmans code executes the QE codes pw.x and ph.x. Subsequently, the MLWF algorithm is carried out through pw2wannier90.x and wannier90.x [pizzi2020]. Steps two through four employ the kcw.x code that also gave the workflow naming [colonna2022].
The QP band structure calculations is given by Eqs. (7) and (8); The orbital resulting in KI-CX results are themselves the KI-CX QP predictions. These results are compared to QP states of KI-PBE and the KS states of the corresponding underlying functionals CX and PBE. KI-PBE calculations are standard in the Koopmans-DFT code. To do band structure calculations with KI-CX it is necessary to explicitly introduce a selection of the CX functional in the ‘QE’ specification, see tutorial Ref. [koopweb2025].
We compute QP states KI-CX (technically, ‘Koopmans DFPT’ [colonna2022] with a CX start) for the h-BN monolayer (BN1) and for bulk h-BN in the experimentally observed AB and AA’ stacking configurations. The structural characters of these h-BN systems are illustrated in Fig. 1 that also identifies the set of relevant high-symmetry $k$-points in the Brillouin zone (BZ).
The BN1 structure (first column) is a monolayer exhibiting the conventional honeycomb lattice, where boron (green) and nitrogen (gray) atoms with valence 3 occupy alternating vertices of the hexagonal network. In-plane $sp^{2}$ bonding between boron and nitrogen atoms provides structural stability and rigidity for this single-sheet system. The choice of in-plane lattice constants for the BN1 system is set by an early experimental structure characterizations of h-BN [hassel1927]: For BN1 we use an hexagonal unit cell with in-plane parameters $a=b=2.504$ Å and a $7\times 7\times 1$ $k$-point grid sampling. In our periodic modeling used for QE studies we set $c=20$ Å; This choice gives sufficient distance between repeated images of the BN1 layer that we can, in practice, ignore spurious inter-layer forces in characterizations of cohesion as well as of QP energies, as discussed in Appendix B.
@@
Bulk AB stacking (second column of Fig. 1) features a graphite-like staggered arrangement of hexagons, where half of the atoms (for example, all of the nitrogens) sit directly on top of a different type of atom in the neighboring layer. For our calculation study we use a $7\times 7\times 3$ $k$-point grid with lattice constants set at an average over measurements on what are interpret [lifesh2003] as experimentally characterized AB structures, $a=b=2.505$ Å and $c=6.661$ Å.
The third column of panels in Fig. 1 show schematics of the $A$ form that is the ground-state of h-BN. Contrary to the AB form, the ground-state configuration stacks the BN layers so that all of the boron (nitrogen) atoms in a given layer sit directly on top of a nitrogen (boron) atom in neighboring layers. For our AA’ studies, we use an hexagonal unit cell with lattice constants parameters $a=b=2.504$ Å and $c=6.66$ Å that represent averages of values extracted by experiments [hassel1927, pease1952, lifesh2003]. Also, we again use a $7\times 7\times 3$ $k$-point sampling.
Additionally, we include a study demonstrating the full vdW impact on the QP band structure, contrasting what we call consistent-KI-CX and consistent-KI-PBE descriptions. That is, we track not only the direct changes arising when we switch from a PBE to a CX starting point in KI-DFT at the same fixed geometry, but also the additional indirect signatures that arise because a GGA PBE is not set up to accurately characterize the bulk h-BN structure [rydberg03p606, rydberg03p606, rohrer11p165423, hyldgaard2020]. The importance of this is discussed in the following section IV. In practice, we complete KI-PBE and KI-CX QP characterizations, having first relaxed the interlayer distance of the AA system using KS-DFT with the PBE and the CX functionals, respectively. Specifically, we find that CX-based and PBE-based structure optimizations yield lattice constants $c=6.448$ Å and $c=8.560$ Å, respectively. The former and latter structure is then used in ensuing KI-CX(CX) and KI-PBE(PBE) QP characterizations.
For the three materials forms at experimental structures, and for the demonstration of indirect vdW fingerprints, we presents analysis of the resulting QP bandstructures, using a shared computational treatment. The Gygi-Baldereschi method [gygi1986] is used to address the $\mathbf{q}\to 0$ divergence of the Coulomb interaction in our periodic systems, both in the initial QE code stages and when setting the screening coefficients $\alpha_{i}$ for the KI-DFT studies, above. The Gygi-Balderashi approach assumes some level of isotropy and its use is more motivated for the bulk BN cases than for the single-sheet BN1 system.
@@ B $p_{z}$ $R_{2}\quantity(r)$
In contrasting predictions of the KI-DFT predictions and KS-DFT approximations for the QP bandstructures we face, in principle, a problem of how to pick a reference energy. Bandstructure characterizations are often reported with QP energies listed relative to the valence band maximum (VBM) but such values are generally affected by both the atom configurations and choice of methodology. Ideally, as the QPs reflect vertical excitations, we would like to list the variation with QP energies defined relative to the true vacuum. However, in a planewave code like QE, we can only get orbital-energy values relative to what is the average electrostatic potential in a given cell, with a given atomic configuration (and choice of pseudopotentials) and for the electron-density variation that is produced in the given DFT method and given choice of XC functional [RMBook].
@@
where $\varepsilon^{KI-CX}_{\texttt{VBM}}$ the VBM of KI-CX and where $M$ corresponds to difference choices of the methodologies: PBE, CX, KI-PBE and KI-CX. The point is that since the atomic and cell geometry are shared (and we use the same pseudopotentials in all studies) the average electrostatic potential can only vary because of differences in the electron-density variation. However, PBE and CX are both widely tested and both documented to provide highly accurate accounts of the electron-density variations (at fixed structure) that must therefore be almost identical in the case of PBE and CX studies. Furthermore, by design of the KI-DFT method, it is also so that the KI-CX and CX yields exactly the same electron density variations (at shared geometries) and the same applies for KI-PBE and PBE [linscott2023]. By the explicit procedure that the QE code uses to set the average electrostatic potentials, Refs. [Tosi1964, RMBook], no differences can therefore arise for the electrostatic reference energy in KI-CX and CX. The same offset brings the here-reported bandstructures in alignment with a QP bandstructure mapping that reflects the true vacuum level.

### v2

@@ II.2 CX in Koopmans DFT
From a theoretical perspective, the eigenstates of KS-DFT, obtained using a regular XC functional, provide a reasonable first approximation for studying charge excitations in [10, 11]. Still, while KS-DFT serves us to accurately predict GS properties, its eigenvalues $\epsilon_{i}^{\text{KS}}$ and orbitals $\phi_{i}^{\text{KS}}(\mathbf{r})$ do not directly describe charged excitations such as the set of vertical ionization energies, i.e., the QPs. This is true even though that all the necessary information to describe a frequency moment of the fully interacting spectral function is, in fact, available by simply combing terms of the KS-DFT energy descriptions with also an evaluation of the kinetic-correlation energy [112, 113, 114]. Even a PBE-based evaluation of those DFT terms yields a value for the moment of the interacting spectral function that closely reflects a full multi-reference configuration integral studies for small molecules [114].
The limitation of using KS-DFT orbitals as a predictor for QPs arises because KS-DFT relies on using (regular, nonhybrid) XC functionals to define an effective local potential contribution [32, 16] to set the orbital structure. Plotting the total energy as a function of the number of electrons then yields a convex curvature [16]. Use of regular XC functionals to thus directly set KS-DFT orbitals implies a lack of the PWL that is, for example, needed for a DFT-type framework to also satisfy a thermodynamics constraints on charge transfers [67, 16].
Additional information, concealed as MBPT input to the XC functional designs, is needed to also recover accurate results for QP levels [24, 16, 21]. Fortunately the information is still formally retained in the MBPT input and physics logic of good XC designs [114]. By switching to Koopmans-compliant DFT and using the KI-DFT framework with a linear-response extensions of an underlying XC functional (like PBE or the CX version of the vdW-DF method), we can leverage more of the underlying MBPT contents [32, 35, 114] of such robust XC designs also for accurate QP predictions [28, 68, 18, 105, 21, 23].
On a formal level, KI-DFT employs a class of orbital-density-dependent functionals (ODDFs) designed to enforce generalized PWL. Specifically, the ODDFs are crafted to enforce that we have a linear energy change with respect to variations in the electron occupation [67, 16, 24]. The CX-based adjustments for KI-DFT, what can also be called the Koopmans functional, is given:
@@
Here $\{\alpha_{i}\}$ are screening parameters that account for electronic relaxation effects and $\rho_{i}$ are orbital-specific density variations $\rho_{i}(\mathbf{r})=f_{i}|\phi_{i}(\mathbf{r})|^{2}$; with $0<f_{i}<1$ [24]. We are using the KI form [26, 49, 28, 29] so that the formal KI-functional specification, $E^{\text{KI-CX}}_{\text{xc}}\left[\rho,\{\rho_{i}\}\right]$, reduces exactly to the underlying CX functional form when there is an integer charge in the unit cell (and hence KI-CX does not change KS-DFT structure predictions obtained by CX). However, for partial orbital occupation, the Koopmans-correction term $\Pi_{i}^{\text{CX}}$ in Eq. (7) restores PWL: It converts the nonlinear dependence of the CX functional on changes in orbital-specific occupations, $\int\rho_{i}(\mathbf{r})d\mathbf{r}$, into a linear form between any two integer occupations. This approach suffices to convert the KS-DFT input (i.e., the choice of the XC) into the KI-DFT tool for making corresponding QP predictions [49, 27, 29]. Further details are provided in Refs. [24, 29].
@@
Although this method rigorously provides both variational orbitals and $\{\alpha_{i}\}$, its application remains challenging due to the requirement of multiple constrained DFT and Koopmans calculations in supercells for periodic systems. To address this limitation, the kcw formulation, discussed in Refs. [115, 28, 27], employs a second-order Taylor expansion of $\Pi_{i}[\rho,\rho_{i}]$ and use of DFPT in primitive cells. This approach assumes coincidence between variational orbitals and maximally localized Wannier functions (MLWFs) $\quantity{w_{i\mathbf{k}}}$) [116], that are computed by use of the same underlying functional. These MLWFs simplify the minimization of $E^{\text{KI-CX}}_{\text{xc}}\left[\rho,\{\rho_{i}\}\right]$.
@@
where $\mathcal{V}^{\text{KI}}_{i\mathbf{0}}$ is the orbital density dependent potential (ODDP) and corresponds to the variation $\delta\sum_{j}\Pi_{j}/\delta\rho_{i}$, and $\hat{h}^{\text{CX}}$ is the effective potential defined by CX functional [14]. The subscript index $\mathbf{0}$ in $\mathcal{V}^{\mathrm{CX}}_{\mathbf{0}i}$ and $\alpha_{\mathbf{0}i}$ refers to a decomposition into a set of independent problems, where a supercell calculation with $\Gamma$-point sampling corresponds to sampling the folded Brillouin zone of the primitive cell [28]. Hamiltonian diagonalization in Eq. (8),
@@
yields so-called canonical orbitals that corresponds to the eigenvectors $\tilde{w}_{\mathbf{k}i}$. The eigenvalues of Eq. (9), $\varepsilon_{\mathbf{k}i}$ define the KI-CX (or KI-DFT) predictions for the QP energies. We note that the canonical orbitals have a Block-wavefunction-like character; They are delocalized as they are linear combination of MLWFs $w_{\mathbf{k}i}$. It has been shown that use of an ODDP can provide a QP approximation to the frequency-dependent contraction of the many-body self-energy term [25]. The implication is that ODDFs have the flexibility to accurately describe both total energies and QP excitations.
@@
Appendix A provides details on a formal representation of a plane-wave and grid-based determination of Eq. (10) using the Roman-Soler spline formulation [96] for (the functional nature [70, 46, 71] of) the nonlocal-correlation kernel $\Phi_{c}^{\text{nl}}[\rho](\mathbf{r},\mathbf{r^{\prime}})$ entering Eq. (6). The formal expressions are already coded in the linear-response module (kcw) of QE thanks to a previous study documenting fingerprints of vdW and nonlocal-correlation interactions on phonon predictions [117].
@@ III Computational details
We compute the QP states by the Koopmans-DFT code package [30, 29], using the KI-DFT versions, and the above-summarized kcw formulation [28]. All those calculations as well corresponding KS-DFT studies are done using norm-conserving PBE-type (scalar-relativistic) pseudopotentials from the PseudoDojo library [118]. That is, we employ dual projectors and generalized norm conservation to accurately reproduce all-electron potential binding and scattering properties. Appendix B presents a discussion of numerical stability for the QP-energy predictions with the wavefunction energy cut off $E_{\rm cut}^{\rm ef}$ (and, for the strictly two-dimensional BN1 system, the height $c$ of the modeling cell). This analysis leads us to use $E_{\rm cut}^{\rm wf}=200$ Ry (400 Ry) for bulk (single-sheet) h-BN, while doubling the grid for the density variations; Exceptions are explicitly stated.
@@
In the initial step, Koopmans code executes the QE codes pw.x and ph.x. Subsequently, the MLWF algorithm is carried out through pw2wannier90.x and wannier90.x [119]. Steps two through four employ the kcw.x code that also gave the workflow naming [28].
The QP band structure calculations is given by Eqs. (7) and (8); The orbital resulting in KI-CX results are themselves the KI-CX QP predictions. These results are compared to QP states of KI-PBE and the KS states of the corresponding underlying functionals CX and PBE. KI-PBE calculations are standard in the Koopmans-DFT code. To do band structure calculations with KI-CX it is necessary to explicitly introduce a selection of the CX functional in the ‘QE’ specification, see tutorial Ref. [30].
We compute QP states KI-CX (technically, ‘Koopmans DFPT’ [28] with a CX start) for the h-BN monolayer (BN1) and for bulk h-BN in the experimentally observed AB and AA’ stacking configurations. The structural characters of these h-BN systems are illustrated in Fig. 1 that also identifies the set of relevant high-symmetry $k$-points in the Brillouin zone (BZ).
The BN1 structure (first column) is a monolayer exhibiting the conventional honeycomb lattice, where boron (green) and nitrogen (gray) atoms with valence 3 occupy alternating vertices of the hexagonal network. In-plane $sp^{2}$ bonding between boron and nitrogen atoms provides structural stability and rigidity for this single-sheet system. The choice of in-plane lattice constants for the BN1 system is set by an early experimental structure characterizations of h-BN [120]: For BN1 we use an hexagonal unit cell with in-plane parameters $a=b=2.504$ Å and a $7\times 7\times 1$ $k$-point grid sampling. In our periodic modeling used for QE studies we set $c=20$ Å; This choice gives sufficient distance between repeated images of the BN1 layer that we can, in practice, ignore spurious inter-layer forces in characterizations of cohesion as well as of QP energies, as discussed in Appendix B.
@@
Bulk AB stacking (second column of Fig. 1) features a graphite-like staggered arrangement of hexagons where half of the atoms (for example, all of the nitrogens) sit directly on top of a different type of atom in neighboring layers. For our AB calculations, we use a $7\times 7\times 3$ $k$-point grid with lattice constants set at an average over measurements on what are interpret [121] as experimentally characterized AB structures, $a=b=2.505$ Å and $c=6.661$ Å.
The third column of panels in Fig. 1 show schematics of the $A$ form that is the ground-state of h-BN. Contrary to the AB form, the ground-state configuration stacks the BN layers so that all of boron (nitrogen) atoms in a given layer sit directly above a nitrogen (boron) atom in neighboring layers. For our AA’ studies, we use an hexagonal unit cell with lattice constants, $a=b=2.504$ Å and $c=6.66$ Å, that represent averages of values extracted by experiments [120, 122, 121]. Also, we again use a $7\times 7\times 3$ $k$-point sampling.
Additionally, we include a study demonstrating the full vdW impact on the QP band structure, contrasting what we call consistent-KI-CX and consistent-KI-PBE descriptions. That is, we track not only the direct changes arising when we switch from a PBE to a CX starting point in KI-DFT at the same fixed geometry, but also the additional indirect signatures that arise because a GGA PBE is not set up to accurately characterize the bulk h-BN structure [123, 123, 124, 35]. The importance of this is discussed in the following section IV. In practice, we complete KI-PBE and KI-CX QP characterizations, having first relaxed the interlayer distance of the AA system using KS-DFT with the PBE and the CX functionals, respectively. Specifically, we find that CX-based and PBE-based structure optimizations yield lattice constants $c=6.448$ Å and $c=8.560$ Å, respectively. The former and latter structure is then used in ensuing KI-CX(CX) and KI-PBE(PBE) QP characterizations.
For the three materials forms at experimental structures, and for the demonstration of indirect vdW fingerprints, we presents analysis of the resulting QP bandstructures, using a shared computational treatment. The Gygi-Baldereschi method [125] is used to address the $\mathbf{q}\to 0$ divergence of the Coulomb interaction in our periodic systems, both in the initial QE code stages and when setting the screening coefficients $\alpha_{i}$ for the KI-DFT studies, above. The Gygi-Balderashi approach assumes some level of isotropy and its use is more motivated for the bulk BN cases than for the single-sheet BN1 system.
@@ B $p_{z}$ $R_{2}\quantity(r)$
In contrasting predictions of the KI-DFT predictions and KS-DFT approximations for the QP bandstructures we face, in principle, a problem of how to pick a reference energy. Bandstructure characterizations are often reported with QP energies listed relative to the valence band maximum (VBM) but such values are generally affected by both the atom configurations and choice of methodology. Ideally, as the QPs reflect vertical excitations, we would like to list the variation with QP energies defined relative to the true vacuum. However, in a planewave code like QE, we can only get orbital-energy values relative to what is the average electrostatic potential in a given cell, with a given atomic configuration (and choice of pseudopotentials) and for the electron-density variation that is produced in the given DFT method and given choice of XC functional [126].
@@
where $\varepsilon^{KI-CX}_{\texttt{VBM}}$ the VBM of KI-CX and where $M$ corresponds to difference choices of the methodologies: PBE, CX, KI-PBE and KI-CX. The point is that since the atomic and cell geometry are shared (and we use the same pseudopotentials in all studies) the average electrostatic potential can only vary because of differences in the electron-density variation. However, PBE and CX are both widely tested and both documented to provide highly accurate accounts of the electron-density variations (at fixed structure) that must therefore be almost identical in the case of PBE and CX studies. Furthermore, by design of the KI-DFT method, it is also so that the KI-CX and CX yields exactly the same electron density variations (at shared geometries) and the same applies for KI-PBE and PBE [29]. By the explicit procedure that the QE code uses to set the average electrostatic potentials, Refs. [127, 126], no differences can therefore arise for the electrostatic reference energy in KI-CX and CX. The same offset brings the here-reported bandstructures in alignment with a QP bandstructure mapping that reflects the true vacuum level.